This problem came from a worksheet given to us. It says: Find 2 values of theta that satisfy cot of theta = -0.2679
Answer by stanbon(75887)

Find 2 values of theta that satisfy cot(theta) = -0.2679
---------------------------------------------------------------
If cot(theta) = -0.2679, tan(theta) = -1/(0.2679)
----------------------
Use your calculator to find tan^-1(-1/0.2679))
I get -75 degrees.
The 2nd angle is -75-180 = -255 degrees
---------------------------------------------
The corresponding positive angles are:
-75 + 360 = 285 degrees
and
-255 + 360 = 105 degrees
===============================
Cheers,
Stan H.